OpenStudy (anonymous):

Solve. |2x + 4| - 5 = 6

OpenStudy (anonymous):

it is x =- 15/2 and x = 7/2

OpenStudy (anonymous):

^

OpenStudy (anonymous):

I would first add 5 to both sides. 2x+4 - 5 = 6 +5 +5 --------------- 2x+4 = 11 now subtract 4 from each side 2x + 4 =11 -4 -4 ------------ 2x = 7 Now divide 2x = 7 /2 /2 -------- x = 7/2

OpenStudy (anonymous):

2x+4-5=6 2x=6-4+5 2x=11-4 2x=7 x=7/2 or x=3 1/2

OpenStudy (anonymous):

check: 2x+4-5=6 2(7/2) +4-5=6 14/2+4-5=6 7+4-5=6 11-5=6 6=6 answer is correct!

OpenStudy (anonymous):

Gee, Thanks RJV

OpenStudy (anonymous):

your welcome..